# Edge Learning for 6G-enabled Internet of Things: A Comprehensive Survey of Vulnerabilities, Datasets, and Defenses

### v1

@@ V Edge Learning for Cyber Security
Edge learning, a variant of machine learning, leverages the power of edge computing—processing data close to its source, thereby reducing latency and bandwidth usage. This technology, when applied to cybersecurity, provides swift, real-time detection and mitigation of cyber threats, thereby boosting the overall resilience of the systems. This section aims to comprehensively discuss the use of edge learning and its application in cybersecurity.
@@ V-A Differential privacy-based solutions
Truex et al. [123] presents a new approach to address the privacy and accuracy trade-offs associated with existing federated learning systems. The authors highlight the vulnerabilities of existing federated learning systems and propose a solution that combines differential privacy and secure multiparty computation. The authors’ approach combines these two techniques that enable the reduction of noise injection as the number of parties increases without sacrificing privacy while maintaining a pre-defined rate of trust. The system is designed to withstand three potential adversaries, including an honest-but-curious aggregator, colluding parties, and outsiders. The system assumes secure communication channels between each party and the aggregator, and the use of the threshold variant of the Paillier encryption scheme. This scheme ensures the privacy of individual messages sent to the aggregator, preventing any set of parties from decrypting ciphertexts. The proposed FL system remains resilient to an inference against potential adversaries who may be users of the service. The experimental results validate the system’s effectiveness and superiority over state-of-the-art solutions, making it a scalable and reliable approach to federated learning. Zhou et al. [124] proposes a privacy-preserving federated learning scheme in fog computing that enables fog nodes to collect Internet-of-Things (IoT) device data and complete the learning task. The proposed scheme addresses the issues of the uneven distribution of data and the large gap of computing power, which affects the efficiency of training and model accuracy in federated learning. The scheme leverages differential privacy to resist data attacks and uses a combination of blinding and Paillier homomorphic encryption to secure model parameters. The proposed scheme is formally verified to guarantee both data security and model security and to resist collusion attacks.
@@
Chen et al. [136] proposes a decentralized global model training protocol, named PPT, that addresses security, privacy preservation, and robustness requirements in the context of Federated Learning in P2P networks. The paper proposes solutions to various security threats and privacy preservation using the symmetric cryptosystem, secure key distribution, and random noise generation. PPT also adopts game theory to resist collusion attacks and has elaborate designs for communication efficiency and dropout-robustness. Extensive experiments show that PPT outperforms Google’s Secure Aggregation and LDP-based FL methods in computation efficiency and prediction performance, but the proposed system is vulnerable to adversarial attacks.
@@ V-B Ensemble Learning-based solution
Ensemble Learning [138] is a machine learning technique that combines multiple individual models to create a more robust and accurate prediction model. When it comes to IoT security, Ensemble Learning can be a valuable tool to enhance the security of IoT devices and networks. Mothukuri et al. [125] proposed a decentralized federated learning approach to enable anomaly detection on Internet of Things (IoT) networks. This approach enables on-device training, thereby eliminating the need to transfer network data to a centralized server, ensuring optimal results in predicting intrusion in IoT networks. The proposed approach utilizes long short-term memory (LSTM) and gated recurrent units (GRUs) neural network models to train the machine learning model on the Modbus network data set. The article also provides an overview of LSTM and GRU networks, their architecture, and real-time implementations. The proposed approach is beneficial in securing data privacy at end devices, achieving optimal results, and being computationally inexpensive.
@@ V-C GAN-based solution
Generative Adversarial Networks (GANs) [139] have been applied successfully in various fields, including computer vision and natural language processing. In recent years, researchers have started exploring their use in IoT security. One potential application of GAN-based solutions is to generate realistic attack scenarios and create diverse datasets for training security models. GANs can also be used to generate synthetic data that mimics real-world sensor data. This can be used to detect anomalies and prevent security breaches. Furthermore, GANs can generate adversarial examples to test the robustness of IoT security systems. Cui et al. [126] presented an approach for decentralized and asynchronous Federated Learning (FL) for Internet of Things (IoT) anomaly detection. The proposed approach uses a modified Generative Adversarial Network (GAN) model called DP-GAN, which has an additional perceptron, the DP identifier (DPI), to generate differential noise, while approximating the raw data to a better degree. The proposed method also integrates blockchain to achieve global aggregation and improve system reliability. The paper demonstrates the effectiveness of the proposed approach in terms of accuracy, privacy protection, and efficient convergence, as shown by the experiments on benchmark datasets.
@@ V-D Fixed-point quantization method-based solution
Fixed-point quantization is a method used in digital signal processing to reduce the storage requirements and computational complexity of numerical data. In IoT security, this technique can be applied to reduce the amount of data transmitted over the network, minimizing the risk of data leakage and unauthorized access. Xu et al. [127] proposes Federated Learning with Minimum Square Quantization Error (FedMSQE), an approach to address the challenges of deploying popular neural network models on Internet of Medical Things (IoMT) devices. The paper focuses on the security of neural network model transmission and reducing the network scale. The proposed method uses fixed-point quantization with stochastic rounding to achieve the smallest quantization error for each individual client in the FL setting. The authors demonstrate through numerical experiments that their proposed algorithm achieves higher accuracy and lower quantization error than other quantization methods, and can be applied to any deep neural network on any single device. FedMSQE improves the security of FL and reduces the transmission cost while maintaining high accuracy.
@@ V-F Homomorphic encryption-based solution
Homomorphic encryption [141] can protect the algorithms and models used in IoT applications, preventing attackers from reverse engineering or tampering with these crucial components. Gao et al. [129] proposed a Homomorphic encryption-based solution, named SVeriFL, that provides a solution to the privacy and security challenges in federated learning, ensuring data integrity and correctness of uploaded parameters and aggregated results. It also allows for consistency verification among multiple participants. The scheme employs a dynamic successive verification mechanism, running throughout the FL training process without impacting the original FL performance. The security analysis and experiments demonstrate the effectiveness of the proposed system, with moderate computational cost.
@@ V-G Blockchain and smart contracts-based solution
The integration of blockchain and smart contracts can offer a robust and decentralized solution for IoT security [142]. By leveraging blockchain technology, data transactions can be recorded and verified by a distributed network of nodes, ensuring that data integrity is maintained and tampering is prevented. Ouyang et al. [130] introduced a privacy framework for FL based on blockchain and smart contracts. The framework offers complete privacy services for off-chain federations that want to use on-chain FL. The proposed framework is implemented and analyzed based on two prevalent blockchain projects, Ethereum and inter-plenary file systems (IPFS). The experiments show that the framework has acceptable collaboration costs and has advantages in privacy, security, and decentralization. Additionally, the framework can enable automatic on-chain identification and autonomous FL of machine clusters made up of IoT devices or distributed participants. Baucas et al. [134] proposes a fog-based IoT platform that uses federated learning and blockchain technology to address the challenges of preserving patient data privacy and improving the security of wearable IoT devices. The platform enforces decentralized servers and resource reallocation to enhance adaptability and sustainability. Federated learning is used to preserve patient data privacy, while blockchain technology provides access control and a cryptographic structure to improve security. The platform’s testbed effectively simulates and evaluates the proposed implementation, and the model accuracy shows the platform’s ability to preserve the integrity of a predictive service.
@@ V-H Dynamic Weighted Aggregation-based solution
The dynamic Weighted Aggregation-based solution is a promising approach for addressing the security challenges associated with IoT. This solution involves integrating multiple security measures and assigning weights to each mechanism based on their importance and effectiveness in a given context. The weights are then dynamically adjusted based on the changing security environment and the behavior of IoT devices. Li et al. [132] introduces DAFL, a dynamic weighted aggregation federated learning system for intrusion detection. The proposed approach reduces the effect of poorly performing local models on the global model during training, which improves the global model’s performance in intrusion detection and reduces communication overhead. The paper presents the system design and key implementation details of DAFL and evaluates its performance against existing approaches. The experimental results demonstrate that DAFL achieves better detection performance with lower communication overhead, but the proposed system does not take into account other factors such as device security or physical attacks.
@@ V-I Software-defined networking
Software-defined networking (SDN) can offer a robust solution for addressing the security risks associated with IoT [143]. SDN-based security solutions allow for fine-grained access control and enforcement policies, ensuring the protection of IoT devices and networks from unauthorized access. Abou El Houda et al. [135] proposes a framework called MiTFed that enables multiple SDN domains to collaboratively build a global intrusion detection model using FL, blockchain, and SDN technologies. The framework achieves high accuracy and efficiency in detecting new and emerging security threats while preserving the privacy of collaborators, making it a promising solution to address security threats in large-scale distributed networks. Further research is needed to assess the scalability of the proposed framework and address potential performance issues due to the use of blockchain technology as well as the dataset used in the performance evaluation is outdated.
@@
TABLE XI: Datasets used as a benchmark experiment in the evaluation of Machine learning vulnerabilities
@@
UCR archive Adversarial attack Black-box adversarial attack FCN and ResNet Centralized Time series classification [159]
@@
LibriSpeech dataset Adversarial attack Black-box adversarial attack DNN Centralized Speech Recognition [163]
@@
UCF-101 dataset Adversarial attack Black-box adversarial attack I3D, CNN+LSTM Centralized Video Recognition [168]
@@ VI Datasets for Edge Learning
Following upon a thorough analysis of recent contributions in the literature, we provide a detailed report on the datasets used by the scientific community for experimenting and evaluating ML techniques on cyber attacks, which is provided in Table XI. Based on the content of each dataset, we classify them into the following seven main categories: 1) IoT-based multi-purpose security, 2) Attack classification, 3) Image classification, 4) Time series classification, 5) Human Activity Recognition (HAR), 6) Sentiment classification, 7) Location awareness, and 8) Text classification, as illustrated in Figure 6.
@@ VI-A IoT-based multi-purpose security
In recent years, addressing IoT security has received a great deal of focus from academia, industry, and governments, mainly motivated by the expanding attack surface from existing threats across the IoT networks. The availability of sophisticated, easy-to-run, and easy-to-find exploit scripts and specialized utility [140]. IoT-based security defenses using AI (specifically anomaly-based approaches) have provided a good line of defense against these threats, and the efforts provided by the security community deserve mention [7]. However, given that one of the most important factors in providing an effective final model is the data. This means that it is safe to say that the final model is well-founded if the data is well-grounded. Although different datasets have been proposed to design, train, and evaluate AI-based security mechanisms for IoT-related environments, the majority of them are either context-specific (e.g., a specific protocol), constrained threat model (limited types and quantities of attacks), or collected within a shallow generation framework (either real or virtual) [56]. Others, however, have been designed as IoT-based multi-purpose security-related datasets such as the Edge-IIoTset dataset.
@@
  • Edge-IIoTset dataset [56]: is a comprehensive and realistic dataset collected within multiple IoT/IIoT environments, intended to be used by ML-based cybersecurity models for training and assessment under both training approaches (centralized and FL). Data are collected from a variety of sources, including notifications, logs, and network traffic. A total of 61 features were selected from 1176 encountered characteristics. In particular, the dataset is produced with a designed IoT/IIoT test bed featuring a broad spectrum of interconnected devices, sensors/actuators, protocols, and multi-layer implementations. The generation framework is designed in a multi-layer approach comprising seven layers, including 1) the IoT/IIoT perception layer where there exist more than 10 types of devices including flame sensors, ultrasonic sensors, and water level detection sensors, 2) the edge layer where various brokers and controllers are implemented 3) the fog layer where various platforms are installed, such as IoT hubs and digital twins 4) the SDN layer 5) the NFV layer 6) the blockchain layer, and 7) the application layer. In addition to general-purpose protocols such as HTTP, the dataset includes a range of IoT and IIoT-specific protocols such as MQTT, CoAP, and Modbus. The dataset includes 14 attack classes, including DoS/DDoS, the man-in-the-middle, malware, and injection attacks. ProvNet-IoT proposed in [144], is a forensic provenance-based scheme for examining IoT-targeted network-level assaults. The system applies progressive provenance within the network to illustrate various events within different attack strategies and generate forensic proofs. The authors employed various community detection algorithms including Label Propagation (LPA). Experimental results on the Edge-IIoTset dataset assessed and demonstrated the potential of ProvNet-IoT to recognize specific selective artifacts in order to generate credible proof in forensic examinations. In the same context (i.e., IoT security), but in a different space, namely, smart health system security, Ghourabi et al. [145] proposed a two-tiered security system. The first one is an IDS for smart medical devices, including the Internet of Medical Things (IoMT), and the second one is a malware detection system for general-purpose devices. The proposed system is built on the basis of an optimized LightGBM and a transformer-based model. Evaluations under the Edge-IIoTset reached accuracies as high as 99%.
@@ VI-B Attack classification
  • KDDCup dataset [182]: since 1999, it has been the most widely used dataset for assessing anomaly detection systems during that decade. It was built on the foundation of roughly $4$ GB of tcpdump data from seven weeks of actual network traffic (collected for the DARPA’98 IDS evaluation program). Within the training dataset, there are about 4.9 million individual connection vectors, with each holding 41 features with labels of either normal or attack. The attack types include DoS, User to Root (U2R), Remote to Local (R2L), and probing attack. Although this dataset is considered somehow outdated, various modern studies continue to use it. For example, the work of Fung et al. [147] proposed a defense mechanism against Sybil-based poisoning attacks, called FoolsGold, which is used for securing FL-based learning. The system proposed places no restriction on the number of expected attackers, nor does it necessitate additional information beyond the learning process, while making minimal assumptions regarding the clients and their data. For the KDDCup dataset, the authors trained a single layer fully-connected softmax for the classification task, using 5 clients performing 5 attacks on the same target class. In the same context, Li et al. [148] proposed LoMar, a Local Malicious Factor defense algorithm for addressing poisoning attacks on FL. The proposed algorithm is based two main phases. The first phase rates each remote client’s model updates by scaling relative distribution across their neighbors using a kernel density estimation technique. In the second phase, a statistically optimal trigger threshold is derived to differentiate between malicious and safe updates. For the KDD Cup dataset, the authors use 2 classes with over two hundred thousand data samples as their source and target label, while setting up three malicious clients with flipped samples. The target label accuracy reported is 99.1%.
@@
  • NSL-KDD dataset [183]: is a proposed dataset to overcome certain issues associated with the KDDCup dataset, including failure to provide a clear definition of the attacks. It consists of two main parts: 1) KDDTrain+ with 125,973 records and 2) KDDTest+ with 22,544 records, which are created from the KDDCup dataset, with four main categories of attacks, namely R2L, Prob, DoS, and U2R. A recent work by Wang et al. [170] proposed IFPA and IUA, two integration-based adversarial generation techniques. These techniques have the potential to conduct DNN-targeted white-box attacks. The IFPA is designed for cases where a specific number of points are to be disturbed, while the IUA is designed for cases where no perturbation point number is requested. The performance of the NSL-KDD dataset achieves an overall acceptable accuracy of 78.96% on the test set.
@@
  • UNSW-NB15 dataset [184]: developed in the Cyber Range Lab at the Australian Cyber Security Centre (ACCS) as a hybrid of real-world and synthetic attack models. It holds over 2.5 million data instances with 49 features mined out of 100 GB of raw traffic. It includes nine types of attack including DoS, shellcode, reconnaissance, generic, exploits, and worms. The work by Zhang et al. [172] proposed SecFedNIDS, a poisoning attacks-robust FL-based NIDS. To effectively shadow poisonous traffic data and avoid its participation in further local training, SecFedNIDS employs a novel detection method for poisonous data based on the similarity of class paths. The authors followed a layer-wise suitability propagation technique to retrieve the classpath from the clean traffic data and transmit it to poisoned clients to assist in discriminating the poisoned data. The results showed that SecFedNIDS increases accuracy in case of poisoning attacks on the UNSW-NB15 dataset by up to 48%.
@@
  • CICIDS2018 dataset [185]: generated by the Canadian Institute for cyber security [185] using the profiles notion containing comprehensive descriptions of attacks and detailed abstract distribution models for applications, protocols, or lower-level network entities. In addition to the Benin profile, the dataset includes different attack types such as DoS/DDoS, botnet, infiltration, web, and brute-force attacks. This dataset has been widely used for evaluating Host and network IDSs in the last few years.
@@
  • DeltaPhish dataset [187]: was gathered from active phishing URLs retrieved online from the PhishTank feed during the period of Oct 2015 to Jan 2016. The authors gathered and verified more than 1000 phishing pages manually. Within each phishing page, they subsequently harvested the related homepage of the hosting domain. Using hyperlink analysis in the HTML code, three to five legitimate pages are collected and manually approved. In total, there are 5,511 distinct web pages included, out of which 1,012 are phishing pages. The authors in [181] proposed a collection of Grey-box attacks on phishing detectors that a phishing adversary can use. The attacks vary according to the knowledge that the adversary has of the specific phishing detector. In addition, the authors introduced an associated defense algorithm, named Protective Operation Chain (POC), which is based on a mix of random feature picking. The feature linkages are exploited to minimize the attacker’s guess of the target’s phishing detector. Experiment evaluations on different public datasets including the DeltaPhish dataset showed that the proposed algorithm is robust to attacks on 13 different classifiers.

### v2

@@ V Edge Learning for Cyber Security
Edge learning, a variant of machine learning, leverages the power of edge computing—processing data close to its source, thereby reducing latency and bandwidth usage. When applied to cybersecurity, this technology provides swift, real-time detection and mitigation of cyber threats, thereby boosting the overall resilience of the systems. This section aims to discuss edge learning and its application in cybersecurity comprehensively.
@@ V-A Differential privacy-based solutions
Truex et al. [92] presents a new approach to address the privacy and accuracy trade-offs associated with existing federated learning systems. The authors highlight the vulnerabilities of existing federated learning systems and propose a solution that combines differential privacy and secure multiparty computation. The authors’ approach combines these two techniques that enable the reduction of noise injection as the number of parties increases without sacrificing privacy while maintaining a pre-defined rate of trust. The system is designed to withstand three potential adversaries: an honest but curious aggregator, colluding parties, and outsiders. The system assumes secure communication channels between each party and the aggregator and uses the threshold variant of the Paillier encryption scheme. This scheme ensures the privacy of individual messages sent to the aggregator, preventing any set of parties from decrypting ciphertexts. The proposed FL system remains resilient to an inference against potential adversaries who may be service users. The experimental results validate the system’s effectiveness and superiority over state-of-the-art solutions, making it a scalable and reliable approach to federated learning. Zhou et al. [93] proposes a privacy-preserving federated learning scheme in fog computing that enables fog nodes to collect Internet-of-Things (IoT) device data and complete the learning task. The proposed scheme addresses the issues of the uneven distribution of data and the large gap of computing power, which affects the efficiency of training and model accuracy in federated learning. The scheme leverages differential privacy to resist data attacks and uses a combination of blinding and Paillier homomorphic encryption to secure model parameters. The proposed scheme is formally verified to guarantee data security and model security, and resist collusion attacks.
@@
Chen et al. [105] proposes a decentralized global model training protocol, named PPT, that addresses security, privacy preservation, and robustness requirements in the context of Federated Learning in P2P networks. The paper proposes solutions to security threats and privacy preservation using the symmetric cryptosystem, secure key distribution, and random noise generation. PPT also adopts game theory to resist collusion attacks and has elaborate communication efficiency and dropout-robustness designs. Extensive experiments show that PPT outperforms Google’s Secure Aggregation and LDP-based FL methods in computation efficiency and prediction performance, but the proposed system is vulnerable to adversarial attacks.
@@ V-B Ensemble Learning-based solution
Ensemble Learning [159] is a machine learning technique that combines multiple individual models to create a more robust and accurate prediction model. When it comes to IoT security, Ensemble Learning can be a valuable tool to enhance the security of IoT devices and networks. Mothukuri et al. [94] proposed a decentralized, federated learning approach to enable anomaly detection on Internet of Things (IoT) networks. This approach enables on-device training, eliminating the need to transfer network data to a centralized server and ensuring optimal results in predicting intrusion in IoT networks. The proposed approach utilizes long short-term memory (LSTM) and gated recurrent units (GRUs) neural network models to train the machine learning model on the Modbus network data set. The article also provides an overview of LSTM and GRU networks, architecture, and real-time implementations. The proposed approach is beneficial in securing data privacy at end devices, achieving optimal results, and being computationally inexpensive.
@@ V-C GAN-based solution
Generative Adversarial Networks (GANs) [160] have been applied successfully in various fields, including computer vision and natural language processing. In recent years, researchers have started exploring their use in IoT security. One potential application of GAN-based solutions is to generate realistic attack scenarios and create diverse datasets for training security models. GANs can also generate synthetic data that mimics real-world sensor data. This can be used to detect anomalies and prevent security breaches. Furthermore, GANs can generate adversarial examples to test the robustness of IoT security systems. Cui et al. [95] presented an approach for decentralized and asynchronous Federated Learning (FL) for Internet of Things (IoT) anomaly detection. The proposed approach uses a modified Generative Adversarial Network (GAN) model called DP-GAN, which has an additional perceptron, the DP identifier (DPI), to generate differential noise, while approximating the raw data to a better degree. The proposed method also integrates blockchain to achieve global aggregation and improve system reliability. The paper demonstrates the effectiveness of the proposed approach in terms of accuracy, privacy protection, and efficient convergence, as shown by the experiments on benchmark datasets.
@@ V-D Fixed-point quantization method-based solution
Fixed-point quantization is a method used in digital signal processing to reduce numerical data’s storage requirements and computational complexity. In IoT security, this technique can reduce the amount of data transmitted over the network, minimizing the risk of data leakage and unauthorized access. Xu et al. [96] proposes Federated Learning with Minimum Square Quantization Error (FedMSQE), an approach to address the challenges of deploying popular neural network models on Internet of Medical Things (IoMT) devices. The paper focuses on the security of neural network model transmission and reducing the network scale. The proposed method uses fixed-point quantization with stochastic rounding to achieve a minor quantization error for each client in the FL setting. The authors demonstrate through numerical experiments that their proposed algorithm achieves higher accuracy and lower quantization error than other quantization methods and can be applied to any deep neural network on any single device. FedMSQE improves the security of FL and reduces the transmission cost while maintaining high accuracy.
@@ V-F Homomorphic encryption-based solution
Homomorphic encryption [162] can protect the algorithms and models used in IoT applications, preventing attackers from reverse engineering or tampering with these crucial components. Gao et al. [98] proposed a Homomorphic encryption-based solution, named SVeriFL, that provides a solution to the privacy and security challenges in federated learning, ensuring data integrity and correctness of uploaded parameters and aggregated results. It also allows for consistency verification among multiple participants. The scheme employs a dynamic successive verification mechanism, running throughout the FL training process without impacting the original FL performance. The security analysis and experiments demonstrate the effectiveness of the proposed system with moderate computational cost.
@@ V-G Blockchain and smart contracts-based solution
Integrating blockchain and smart contracts can offer a robust and decentralized solution for IoT security [163]. By leveraging blockchain technology, data transactions can be recorded and verified by a distributed network of nodes, ensuring that data integrity is maintained and tampering is prevented. Ouyang et al. [99] introduced a privacy framework for FL based on blockchain and smart contracts. The framework offers complete privacy services for off-chain federations that want to use on-chain FL. The proposed framework is implemented and analyzed based on two prevalent blockchain projects, Ethereum and inter-plenary file systems (IPFS). The experiments show that the framework has acceptable collaboration costs and has advantages in privacy, security, and decentralization. Additionally, the framework can enable automatic on-chain identification and autonomous FL of machine clusters of IoT devices or distributed participants. Baucas et al. [103] proposes a fog-based IoT platform that uses federated learning and blockchain technology to address the challenges of preserving patient data privacy and improving the security of wearable IoT devices. The platform enforces decentralized servers and resource reallocation to enhance adaptability and sustainability. Federated learning preserves patient data privacy, while blockchain technology provides access control and a cryptographic structure to improve security. The platform’s testbed effectively simulates and evaluates the proposed implementation, and the model accuracy shows the platform’s ability to preserve the integrity of a predictive service.
@@ V-H Dynamic Weighted Aggregation-based solution
The dynamic Weighted Aggregation-based solution is a promising approach for addressing the security challenges associated with IoT. This solution integrates multiple security measures and assigns weights to each mechanism based on their importance and effectiveness in a given context. The weights are then dynamically adjusted based on the changing security environment and the behavior of IoT devices. Li et al. [101] introduces DAFL, a dynamic weighted aggregation federated learning system for intrusion detection. The proposed approach reduces the effect of poorly performing local models on the global model during training, which improves the global model’s performance in intrusion detection and reduces communication overhead. The paper presents the system design and key implementation details of DAFL and evaluates its performance against existing approaches. The experimental results demonstrate that DAFL achieves better detection performance with lower communication overhead. Still, the proposed system does not consider other factors such as device security or physical attacks.
@@ V-I Software-defined networking
Software-defined networking (SDN) can offer a robust solution for addressing the security risks associated with IoT [164]. SDN-based security solutions allow for fine-grained access control and enforcement policies, protecting IoT devices and networks from unauthorized access. Abou El Houda et al. [104] proposes a framework called MiTFed that enables multiple SDN domains to collaboratively build a global intrusion detection model using FL, blockchain, and SDN technologies. The framework achieves high accuracy and efficiency in detecting new and emerging security threats while preserving the privacy of collaborators, making it a promising solution to address security threats in large-scale distributed networks. Further research is needed to assess the scalability of the proposed framework and address potential performance issues due to the use of blockchain technology. The dataset used in the performance evaluation is outdated.
@@
TABLE XII: Datasets used as a benchmark experiment in the evaluation of Machine learning vulnerabilities
@@
UCR archive Adversarial attack blue-box adversarial attack FCN and ResNet Centralized Time series classification [180]
@@
LibriSpeech dataset Adversarial attack blue-box adversarial attack DNN Centralized Speech Recognition [184]
@@ VI Datasets for Edge Learning
Following upon a thorough analysis of recent contributions in the literature, we provide a detailed report on the datasets used by the scientific community for experimenting and evaluating ML techniques on cyber attacks, which is provided in Table XII. Based on the content of each dataset, we classify them into the following nine main categories: 1) IoT-based multi-purpose security, 2) Attack classification, 3) Electricity theft detection, 4) Image classification, 5) Time series classification, 6) Human Activity Recognition, 7) Sentiment classification, 8) Location awareness, and 9) Text classification, as illustrated in Figure 8.
@@ VI-A IoT-based multi-purpose security
In recent years, addressing IoT security has received much focus from academia, industry, and governments, mainly motivated by the expanding attack surface from existing threats across the IoT networks. The availability of sophisticated, easy-to-run, and easy-to-find exploit scripts and specialized utility [161]. IoT-based security defenses using AI (specifically anomaly-based approaches) have provided a good line of defense against these threats, and the efforts provided by the security community deserve mention [10]. However, given that one of the most critical factors in providing an effective final model is the data. This means that it is safe to say that the final model is well-founded if the data is well-grounded. Although different datasets have been proposed to design, train, and evaluate AI-based security mechanisms for IoT-related environments, the majority of them are either context-specific (e.g., a specific protocol), constrained threat model (limited types and quantities of attacks), or collected within a shallow generation framework (either real or virtual) [123]. Others, however, have been designed as IoT-based multi-purpose security-related datasets such as the Edge-IIoTset dataset.
@@
  • Edge-IIoTset dataset [123]: is a comprehensive and realistic dataset collected within multiple IoT/IIoT environments, intended to be used by ML-based cybersecurity models for training and assessment under both training approaches (centralized and FL). Data are collected from a variety of sources, including notifications, logs, and network traffic. A total of 61 features were selected from 1176 encountered characteristics. In particular, the dataset is produced with a designed IoT/IIoT test bed featuring a broad spectrum of interconnected devices, sensors/actuators, protocols, and multi-layer implementations. The generation framework is designed in a multi-layer approach comprising seven layers, including 1) the IoT/IIoT perception layer where there exist more than 10 types of devices including flame sensors, ultrasonic sensors, and water level detection sensors, 2) the edge layer where various brokers and controllers are implemented 3) the fog layer where various platforms are installed, such as IoT hubs and digital twins 4) the SDN layer 5) the NFV layer 6) the blockchain layer, and 7) the application layer. In addition to general-purpose protocols such as HTTP, the dataset includes a range of IoT and IIoT-specific protocols such as MQTT, CoAP, and Modbus. The dataset includes 14 attack classes: DoS/DDoS, man-in-the-middle, malware, and injection attacks. ProvNet-IoT proposed in [165], is a forensic provenance-based scheme for examining IoT-targeted network-level assaults. The system applies progressive provenance within the network to illustrate various events within different attack strategies and generate forensic proofs. The authors employed various community detection algorithms including Label Propagation (LPA). Experimental results on the Edge-IIoTset dataset assessed and demonstrated the potential of ProvNet-IoT to recognize specific selective artifacts in order to generate credible proof in forensic examinations. In the same context (i.e., IoT security), but in a different space, namely, smart health system security, Ghourabi et al. [166] proposed a two-tiered security system. The first is an IDS for smart medical devices, including the Internet of Medical Things (IoMT), and the second is a malware detection system for general-purpose devices. The proposed system is built using an optimized LightGBM and a transformer-based model. Evaluations under the Edge-IIoTset reached accuracies as high as 99%.
@@ VI-B Attack classification
  • KDDCup dataset [202]: since 1999, it has been the most widely used dataset for assessing anomaly detection systems during that decade. It was built on the foundation of roughly $4$ GB of tcpdump data from seven weeks of actual network traffic (collected for the DARPA’98 IDS evaluation program). Within the training dataset, there are about 4.9 million individual connection vectors, with each holding 41 features with labels of either normal or attack. The attack types include DoS, User to Root (U2R), Remote to Local (R2L), and probing attack. Although this dataset is considered outdated, various modern studies continue using it. For example, the work of Fung et al. [168] proposed a defense mechanism against Sybil-based poisoning attacks called FoolsGold, which is used for securing FL-based learning. The proposed system places no restriction on the number of expected attackers, nor does it necessitate additional information beyond the learning process while making minimal assumptions regarding the clients and their data. For the KDDCup dataset, the authors trained a single layer fully-connected softmax for the classification task, using 5 clients performing 5 attacks on the same target class. In the same context, Li et al. [169] proposed LoMar, a Local Malicious Factor defense algorithm for addressing poisoning attacks on FL. The proposed algorithm is based two main phases. The first phase rates each remote client’s model updates by scaling relative distribution across their neighbors using a kernel density estimation technique. In the second phase, a statistically optimal trigger threshold is derived to differentiate between malicious and safe updates. For the KDD Cup dataset, the authors use 2 classes with over two hundred thousand data samples as their source and target label, while setting up three malicious clients with flipped samples. The target label accuracy reported is 99.1%.
@@
  • NSL-KDD dataset [203]: is a proposed dataset to overcome specific issues associated with the KDDCup dataset, including failure to define the attacks clearly. It consists of two main parts: 1) KDDTrain+ with 125,973 records and 2) KDDTest+ with 22,544 records, which are created from the KDDCup dataset, with four main categories of attacks, namely R2L, Prob, DoS, and U2R. A recent work by Wang et al. [190] proposed IFPA and IUA, two integration-based adversarial generation techniques. These techniques have the potential to conduct DNN-targeted white-box attacks. The IFPA is designed for cases where a specific number of points are to be disturbed, while the IUA is designed for cases where no perturbation point number is requested. The performance of the NSL-KDD dataset achieves an overall acceptable accuracy of 78.96% on the test set.
@@
  • UNSW-NB15 dataset [204]: developed in the Cyber Range Lab at the Australian Cyber Security Centre (ACCS) as a hybrid of real-world and synthetic attack models. It holds over 2.5 million data instances with 49 features mined from 100 GB of raw traffic. It includes nine attack types: DoS, shellcode, reconnaissance, generic, exploits, and worms. The work by Zhang et al. [192] proposed SecFedNIDS, a poisoning attacks-robust FL-based NIDS. To effectively shadow poisonous traffic data and avoid its participation in further local training, SecFedNIDS employs a novel detection method for poisonous data based on the similarity of class paths. The authors followed a layer-wise suitability propagation technique to retrieve the classpath from the clean traffic data and transmit it to poisoned clients to assist in discriminating the poisoned data. The results showed that SecFedNIDS increases accuracy in case of poisoning attacks on the UNSW-NB15 dataset by up to 48%.
@@
  • CICIDS2018 dataset [205]: generated by the Canadian Institute for cyber security [205] using the profiles notion containing comprehensive descriptions of attacks and detailed abstract distribution models for applications, protocols, or lower-level network entities. In addition to the Benin profile, the dataset includes different attack types such as DoS/DDoS, botnet, infiltration, web, and brute-force attacks. This dataset has been widely used for evaluating Host and network IDSs in recent years.
@@
  • DeltaPhish dataset [207]: was gathered from active phishing URLs retrieved online from the PhishTank feed from Oct 2015 to Jan 2016. The authors gathered and verified more than 1000 phishing pages manually. Within each phishing page, they subsequently harvested the related homepage of the hosting domain. Three to five legitimate pages are collected and manually approved using hyperlink analysis in the HTML code. 5,511 distinct web pages are included, out of which 1,012 are phishing pages. The authors in [201] proposed a collection of Grey-box attacks on phishing detectors that a phishing adversary can use. The attacks vary according to the adversary who knows the specific phishing detector. In addition, the authors introduced an associated defense algorithm named Protective Operation Chain (POC), which is based on a mix of random feature picking. The feature linkages are exploited to minimize the attacker’s guess of the target’s phishing detector. Experiment evaluations on different public datasets including the DeltaPhish dataset showed that the proposed algorithm is robust to attacks on 13 different classifiers.